OpenStudy (anonymous):

5^(x+1) = 6^(2x) solve for x and I'm pretty sure you have to use log or ln.

sam (.sam.):

yes

OpenStudy (anonymous):

if you can please help me work this out?

sam (.sam.):

take the natural log of both sides of the equation to remove the variable from the exponent ln(5^(x+1))=ln(6^(2x)) xln(5)+ln(5)=ln(6^(2x)) Simplify the result xln(5)+ln(5)=ln((36)^(x)) xln(5)+ln(5)-xln(36)=0 Factoring x(ln(5)-ln(36))=-ln(5) solve for x from here

OpenStudy (anonymous):

Alright thank you so much @.Sam.

sam (.sam.):

welcome